Sometimes a person cannot clearly see objects close up or far away. To correct this type of vision, bifocals are often used. The top half of the lens is used to view distant objects and the bottom half of the lens is used to view objects close to the eye. A person can clearly see objects only if they are located between 35 cm and 200 cm away from her eyes. Bifocal lenses are used to correct her vision.

What power lens (in diopters) should be used in the bottom half of the lens to allow her to clearly see objects 25 cm away?
Answer #1

Using lens formula,

1/f = 1/v + 1/u

1/f = 1/25 - 1/35

f = 87.5 cm

Power, P = 100/87.5 = 1.14 D

Comment in case any doubt please rate my answer ....